FlavR Co stock has a beta of 2.10, the current risk-free rate is 2.10 percent, and the expected return on the market is 9.10 percent. What is FlavR Co's cost of equity?
Answer #1

Cost of Equity = Risk Free Rate + Beta * (Market Return - Risk Free Rate)

= 2.10 + 2.10 * (9.10 - 2.10)

= 2.10 + 2.10 * 7

= 2.10 + 14.70

Cost of Equity = 16.80%